# MSTN: Fast and Efficient Multivariate Time Series Model

### v1

@@ 2.3 Time-Series Classification
Table 3 presents the classification results across 10 UEA standard benchmark datasets. The proposed MSTN framework is compared against SOTA time-series classification methods, including GPT-2(6) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], FEDformer [zhou2022fedformer], and LightTS [zhang2022morefastmultivariatetime]. MSTN demonstrates competitive performance in multiple datasets. MSTN-Transformer achieves perfect classification (100%) on JapaneseVowels and near-perfect performance on SpokenArabic (99.32%), while MSTN-BiLSTM achieves improved accuracy on Heartbeat (81.49%), JapaneseVowels (100%) and PEMS-SF (96.59%).
@@
Rodeg.[DARUS33012023] Sim. Hum. DE. BiLSTM 99.20 0.992 0.993 0.991 4.39 0.81 99.18 5.07 98.64 4.80 91.00[RF,GB][Rodegast2024]
@@
Boubez.[BOUBEZOUL2020105577] Real. Hum. FR. BiLSTM 93.75 0.945 0.940 0.950 4.16 2.79 93.00 4.20 91.37 4.68 91.59[DT][fi15100333]
@@ 2.4 Generalizability Study
As summarized in Table 3, the proposed MSTN framework demonstrates remarkable generalization in seven international benchmark data sets such as human safety, activity recognition, animal welfare, and mechanical prognostics. Without domain-specific tuning, MSTN shows improved performance over the latest time series SOTA models, including TimesNet [wu2023timesnettemporal2dvariationmodeling] and PatchTST [nie2023timeseriesworth64] in both accuracy and computational efficiency.
MSTN achieves competitive performance across fundamentally different domains, demonstrating its capabilities as a universal temporal model. In human safety applications, MSTN-Transformer reaches 99.53% accuracy on German collision prediction data[DARUS33012023], substantially outperforming prior work (91.0% RF,GB [Rodegast2024]). For activity recognition, MSTN-BiLSTM achieves 96.59% accuracy on the UCI-HAR dataset[har240], representing a 13.2% improvement over TimesNet. The framework also excels in real-world fall detection with 93.75% accuracy, surpassing dedicated DT-based methods. MSTN-BiLSTM also shows strong performance in animal welfare[DISSANAYAKE2025111462] assessment with 93.00% accuracy on calf behavior recognition. Furthermore, MSTN-Transformer demonstrates exceptional capability in mechanical prognostics, achieving 11.26 RMSE on the NASA Turbofan dataset[NasaTurboFan] and significantly outperforming all baseline predictors in remaining useful life estimation.
MSTN delivers this performance with good computational efficiency. For instance, on the Rodegast [DARUS33012023] Germany data set, the MSTN-Transformer variant achieves 0.155 ms/sample inference latency 33× faster than TimesNet and 31.0× faster than PatchTST while maintaining compact model size. This efficiency enables deployment on resource-constrained edge devices for real-time safety applications. MSTN achieves consistent SOTA performance across fundamentally different domains from behavioral analytics to mechanical systems, without any architectural modifications.
@@ 2.5 Ablation Study
Table 4 presents an ablation study evaluating the effectiveness of the proposed MSTN-Transformer components across six datasets (ECL, Weather, ETTm1, ETTh2, Traffic, and ILI) and the MSTN-BiLSTM architecture components on the Weather, ETTm2, and Traffic datasets. To verify the contribution of each component, we compare the full MSTN-Transformer model against five variants: w/o CNN, which removes the convolutional pathway; w/o Transformer, which eliminates the core Transformer module; w/o SE, which removes the squeeze-and-excitation blocks; w/o MHTA, which excludes the Multi-Head Temporal Attention; and w/o Gated F., which removes the gated fusion mechanism. In MSTN-BiLSTM architecture, w/o BiLSTM denotes the variant of MSTN-BiLSTM without the BiLSTM layer to assess its importance in temporal sequence modeling. These results validate our architectural design choices and demonstrate the complementary nature of the integrated components.
Table 4: Ablation study of MSTN-Transformer and MSTN-BiLSTM. $H\in\{96,192,336,720\}$ and for ILI $H\in\{24,36,48,60\}$.
Methods MSTN Full w/o CNN w/o Transf. w/o SE w/o MHTA w/o Gated F.
@@
Methods MSTN Full w/o CNN w/o BiLSTM w/o SE w/o MHTA w/o Gated F.
@@
96 0.085 0.202 0.086 0.202 0.102 0.222 0.085 0.202 0.087 0.204 0.087 0.205
192 0.087 0.205 0.088 0.207 0.101 0.221 0.087 0.205 0.088 0.208 0.088 0.209
336 0.086 0.204 0.086 0.205 0.105 0.227 0.089 0.213 0.090 0.214 0.090 0.215
720 0.087 0.207 0.089 0.210 0.102 0.224 0.092 0.218 0.091 0.217 0.093 0.218
@@ 2.6 Model Complexity and Edge-AI Deployability
The MSTN-Transformer and MSTN-BiLSTM architectures exemplify a fundamental design trade-off in temporal modeling. Both architectures integrate efficient parallel components CNN pathways ($\mathcal{O}(T)$) with either Transformer ($\mathcal{O}(T^{2})$) or BiLSTM ($\mathcal{O}(T)$) mechanisms, combined with $\mathcal{O}(T)$ fusion operations and finalized by a multi-head temporal attention layer of $\mathcal{O}(T^{2})$ complexity. This quadratic component dictates the overall asymptotic complexity of $\mathcal{O}(T^{2})$. However, this design preserves robust temporal modeling capabilities and achieves improved performance and practical efficiency compared to pure $\mathcal{O}(T^{2})$ Transformer-based models. Linear models such as DLinear and TS-Mixer achieve optimal $\mathcal{O}(T)$ complexity, while spectral and frequency-based approaches such as FEDformer and TimesNet attain $\mathcal{O}(T)$ and $\mathcal{O}(T\log T)$ complexity, respectively. The recently proposed MTST further reduces the cost to $\mathcal{O}(J_{b_{n}^{\prime}}^{2})$ via patching. Although efficient, these methods often trade off capacity for unconstrained long-range dependency modeling.
Following the results and inference time metrics reported in SOFTS [han2024softsefficientmultivariatetime] on the Traffic data set (lookback window $L=96$, horizon $H=720$), we compare our results with their metrics reported in Table 5. The MSTN-Transformer achieves improved performance with 0.020 MSE and 0.72 ms inference time, representing 22.4× accuracy improvement and 34.7× speedup over SOFTS, achieving a favorable balance of accuracy and speed. This improved efficiency enables substantially reduced memory usage and inference latency, positioning MSTN as a practical solution for deployment on constrained hardware platforms.
Table 5: A: Performance Analysis (Traffic Dataset, $L=96$, $H=720$). B: MSTN-Trans. Inference Time Component Contributions. Red/Blue: First/Second ranks.
MSE Perf. Gap Inf. Time (ms) Speed Gap
A: SOTA MSE and Efficiency
@@
B: MSTN-Trans. Component Contributions Ablation Study (Traffic Dataset, $H=720$)
Variant Size (MB) Time (ms) Speed Gap Role
@@
[…]
As shown in Table 3, MSTN-Transformer delivers competitive performance across domains, attaining 99.53% accuracy on Rodegast [DARUS33012023] while maintaining efficient inference at 0.155 ms per sample with only 0.54 MB model size. This represents up to 33× speedup over TimesNet (5.07 ms) and 31.0× over PatchTST (4.80 ms), demonstrating favorable performance-efficiency trade-offs across diverse datasets. The combination of compact model sizes (0.54–7.14 MB), improved accuracy, and millisecond-level latency positions MSTN as an ideal solution for complex multivariate time series analysis tasks in edge computing environments. These results demonstrate that capable temporal models can achieve practical deployability in edge environments. The small model footprint satisfies memory constraints, millisecond-level latency enables real-time operation, and high accuracy ensures reliability. This collectively addresses core edge-AI challenges for applications like safety monitoring, predictive maintenance, and activity recognition.
3 Discussion
This work introduces MSTN, a DL architecture founded on a hierarchical multi-scale modeling principle. It integrates three core components: a multi-scale convolutional encoder that constructs a hierarchical feature pyramid; sequential modeling component for global dependency modeling, empirically validated with BiLSTM and Transformer variants; and a gated fusion mechanism augmented with SE and MHTA to enable dynamic, context-aware feature integration. Existing methods often rely on fixed-scale structural priors (e.g., patch-based tokenization, fixed frequency transformations, or frozen backbone architectures), forcing a compromise between scale diversity and inference speed. In contrast, our approach efficiently integrates convolutional local pattern extraction with global dependency modeling via an adaptive gated fusion mechanism. This enables our model to capture multi-scale temporal features without compromising computational efficiency, ultimately achieving state-of-the-art performance.
Comprehensive evaluations demonstrate that the MSTN-Transformer establishes a new SOTA across forecasting, imputation, classification, and generalizability. It achieves a 22.4× improvement in accuracy over the SOFTS model while being 34.7× faster in inference, all within a compact 2.87 MB footprint, while MSTN BiLSTM shows competitive performance. This dual advancement in both performance and efficiency is a key differentiator from prior work. The model also exhibits exceptional cross-domain generalization, with its small model size (0.54–7.14 MB) making it suitable for edge deployment. Ablation studies confirm that each component is critical, with the transformer core providing a foundational capability and the multi-scale and fusion modules delivering significant competitive gains.
Future work will explore cross-variate attention mechanisms, and large-scale pre-training to further enhance MSTN’s capabilities as a foundation model for time series understanding. Our work paves the way for efficient multi-scale temporal modeling under resource constraints, enabling real-time forecasting applications across diverse domains from healthcare to industrial monitoring.

### v2

@@ 2.4 Time-Series Classification
Rodegast[DARUS33012023] Sim. Hum. DE. BiLSTM 99.20 0.992 0.993 0.991 4.39 0.81 99.18 5.07 98.64 4.80 91.00[RF,GB][Rodegast2024]
@@
Boubezoul[BOUBEZOUL2020105577] Real. Hum. FR. BiLSTM 93.75 0.945 0.940 0.950 4.16 2.79 93.00 4.20 91.37 4.68 91.59[DT][fi15100333]
@@
Classification involves categorizing entire time series sequences. It is used to identify which type of event or pattern the sequence represents. For example, this could be the classification of a segment of sensor data as walking, running, falling, or collision activity, which is fundamental for automated monitoring and diagnosis systems. Performance is evaluated using the standard metric of classification accuracy.
Table 3 presents the classification results across 10 UEA standard benchmark datasets. The proposed MSTN framework is compared against SOTA time-series classification methods, including TimesNet [wu2023timesnettemporal2dvariationmodeling], GPT-2(6) [zhou2023fitsallpowergeneraltime], FEDformer [zhou2022fedformer], and LightTS [zhang2022morefastmultivariatetime]. MSTN demonstrates competitive performance in multiple datasets. MSTN-Transformer achieves perfect classification (100%) on JapaneseVowels and near-perfect performance on SpokenArabic (99.32%), while MSTN-BiLSTM achieves improved accuracy on Heartbeat (81.49%), JapaneseVowels (100%) and PEMS-SF (96.59%).
@@ 2.5 Generalizability Study
A robust model should not be a specialist in just one area; it should adapt to new tasks and data distributions across different data and domains. This experiment evaluates the ability of the MSTN to transfer its knowledge across diverse domains, from human activity recognition to mechanical fault prediction. As summarized in Table 3, the proposed MSTN framework demonstrates generalization in seven international benchmark data sets such as human safety, activity recognition, animal welfare, and mechanical prognostics. Without domain-specific tuning, MSTN shows improved performance over the latest time series SOTA models, including TimesNet [wu2023timesnettemporal2dvariationmodeling] and PatchTST [nie2023timeseriesworth64].
MSTN achieves competitive performance across fundamentally different domains, demonstrating its capabilities as a universal temporal model. In human safety applications, MSTN-Transformer reaches 99.53% accuracy on German collision prediction data [DARUS33012023], substantially outperforming prior work (91.0% RF,GB [Rodegast2024]). For activity recognition, MSTN-BiLSTM achieves 96.59% accuracy on the UCI-HAR dataset [har240], representing a 13.2% improvement over TimesNet. The framework also excels in real-world fall detection with 93.75% accuracy, surpassing dedicated DT-based methods. MSTN-BiLSTM also shows strong performance in animal welfare evaluation [DISSANAYAKE2025111462], assessment with 93.00% accuracy on calf behavior recognition. Furthermore, MSTN-Transformer demonstrates better capability in mechanical prognostics, achieving 11.26 RMSE on the NASA Turbofan data set [NasaTurboFan] and significantly outperforming all baseline predictors in the estimate of the remaining useful life.
[…]
@@ 2.6 Ablation Study
Table 4: Ablation study of MSTN-Transformer and MSTN-BiLSTM. $H\in\{96,192,336,720\}$ and for ILI $H\in\{24,36,48,60\}$. Red and blue indicate first and second best results, respectively.
Methods MSTN Full w/o CNN w/o Transf. w/o SE w/o MHA w/o SGF
@@
Methods MSTN Full w/o CNN w/o BiLSTM w/o SE w/o MHA w/o SGF
@@
96 0.085 0.202 0.089 0.208 0.090 0.209 0.086 0.203 0.087 0.204 0.087 0.205
192 0.087 0.205 0.092 0.218 0.101 0.221 0.088 0.207 0.088 0.208 0.088 0.209
336 0.086 0.204 0.095 0.224 0.105 0.227 0.089 0.213 0.090 0.214 0.090 0.215
720 0.087 0.207 0.098 0.227 0.107 0.229 0.091 0.218 0.091 0.217 0.093 0.218
An ablation study systematically removes individual components from the full model to evaluate their contributions. Table 4 presents the results for MSTN-Transformer across six diverse benchmarks ECL (electricity), Weather, ETTm1, ETTh2, Traffic, ILI and for MSTN-BiLSTM on three datasets: Weather, ETTm2, and Traffic. For MSTN-Transformer, five variants are tested: w/o CNN, w/o Transformer, w/o SE, w/o MHA, and w/o SGF. While In MSTN-BiLSTM architecture, w/o BiLSTM denotes the variant of MSTN-BiLSTM without the BiLSTM layer.
In the MSTN-Transformer variant, the results consistently show that the full MSTN-Transformer model achieves the best performance (ranked first in MSE/MAE) across nearly all prediction horizons and datasets. Removing the core Transformer module (w/o Transformer) leads to the most severe performance drop in most cases. For instance, on ETTh2 (a challenging dataset with high volatility), the MSE increases substantially from 0.297 (full model) to 0.385 at horizon 96 when the Transformer is ablated, highlighting its critical role in modeling complex temporal patterns. Similarly, the CNN pathway (w/o CNN) is highly important, especially on datasets like Traffic and ECL where local features are crucial; its removal causes noticeable degradation (e.g., Traffic H=96 MSE rises from 0.017 to 0.020). The SE blocks (w/o SE) also contribute significantly, often yielding the second-best or competitive results when present. For example, on Weather, the SE-ablated variant remains close to the full model, but on ETTm1, its removal increases MSE from 0.127 to 0.140 at H=336. The MHA (w/o MHA) and self-gated fusion (w/o SGF) modules further enhance robustness, with their absence leading to consistent, though sometimes smaller, declines across datasets.
For MSTN-BiLSTM, the full model outperforms all ablated variants. Removing the BiLSTM layer (w/o BiLSTM) results in significant deterioration, particularly on complex datasets like ETTm2 where MSE at H=720 jumps from 0.576 to 0.671. This underscores the importance of recurrent modeling in capturing sequential dependencies. In summary, the ablation study validates that each architectural component is essential. The Transformer or BiLSTM, convolutional pathway, and attention-based mechanisms operate synergistically, with the full integrated model delivering the best overall forecasting accuracy across diverse temporal datasets.
@@ 2.7 Model Complexity and Edge-AI Deployability
2.7.1 Complexity Analysis
In this analysis, we denote the length of the lookback window as $L$, the number of variate features (channels) as $C$, and the task-specific output dimension as $K$. We use $C$ for channels to align with notation in common literature such as SOFTS [han2024softsefficientmultivariatetime], where $K$ is often referred to as the forecasting horizon ($H$). We simplify the asymptotic complexity expressions by omitting the model dimension $d$, assuming $d\approx C$.
The MSTN framework addresses the fundamental trade-off in temporal modeling between representational capacity and computational efficiency. Although the proposed variants exhibit distinct encoding complexities with MSTN-BiLSTM maintaining strict linearity with respect to sequence length and MSTN-Transformer incurring a quadratic cost. They both converge on a unified efficiency strategy: ETA. This is achieved because both the parallel CNN and sequence modeling pathways utilize immediate temporal pooling to eliminate sequence-length dependence in subsequent layers. Specifically, both the parallel CNN and sequence modeling pathways (Transformer/BiLSTM) independently condense their temporal features via Global Average Pooling and Sequence Mean Pooling, respectively (reducing the sequence dimension from $L$ to $1$). Following concatenation, the subsequent SGF, SE-Block, and MHA refinement stages operate on static feature vectors. This strategic design ensures that the refinement pipeline scales as $\mathcal{O}(C^{2})$ with respect to channels but remains constant-time ($\mathcal{O}(1)$) with respect to the input length $L$, effectively isolating the heavy temporal computation to the initial stage.
Component-Wise Analysis of MSTN-Transformer: The complexity is dominated by the sequence encoding stage. First, the Multi-Scale CNN extracts local features with linear complexity $\mathcal{O}(CL)$. Second, the Transformer Encoder applies self-attention; as this requires computing an $L\times L$ attention matrix, it introduces the dominant quadratic term $\mathcal{O}(CL^{2})$. Following this, Temporal Pooling aggregates the sequence with linear cost $\mathcal{O}(CL)$. Crucially, the Refinement stage operates on the pooled vector, scaling with channel interactions as $\mathcal{O}(C^{2})$. Finally, the Projection Head maps features to the output with complexity $\mathcal{O}(CK)$. Summing these components, the total asymptotic complexity is $\mathcal{O}(CL^{2}+C^{2}+CK)$, which scales quadratically with sequence length $L$.
Component-Wise Analysis of MSTN-BiLSTM: In contrast, the MSTN-BiLSTM variant achieves strict linear efficiency with respect to sequence length. The BiLSTM Encoder processes the sequence recurrently without generating a global attention map, resulting in a linear complexity of $\mathcal{O}(CL)$. Similar to the Transformer variant, Temporal Pooling ($\mathcal{O}(CL)$) condenses the time dimension. The Refinement stage contributes $\mathcal{O}(C^{2})$, and the Projection Head contributes $\mathcal{O}(CK)$. Since the encoder avoids quadratic operations in $L$, the total start-to-end complexity remains highly efficient at $\mathcal{O}(CL+C^{2}+CK)$, which scales linearly with $L$.
Table 5: Comprehensive evaluation of MSTN across diverse time-series benchmarks. Inference time of MSTN compared to models including SOFTS, iTransformer, TimesNet and PatchTST. Red/Blue: First/Second ranks.
(a) Comparison of MSTN with the MSE and inference time (I.T) metrics reported in SOFTS [han2024softsefficientmultivariatetime] on the Traffic dataset ($L=96$, $H=720$).
MSE Perf. Gap I.T.(ms) Speed Gap
@@ 2.7.1 Complexity Analysis
(b) Comprehensive evaluation of MSTN across diverse time-series forecasting benchmarks [wu2022autoformer].
Dataset Horizon Size (MB) I.T. (ms)
Traffic 96,192,336,720 2.87 0.72
ETTh2 96,192,336,720 3.65 0.91
ETTm2 96,192,336,720 2.92 0.85
Weather 96,192,336,720 2.85 0.70
Exchange 96,192,336,720 2.96 0.80
ILI 24,36,48,60 2.51 0.63
(c) Inference time (I.T., in ms) of MSTN variants compared to TimesNet and PatchTST across seven international datasets.
Dataset MSTN Size (MB) I.T. TimesNet I.T. PatchTST I.T.
Rodegast [DARUS33012023] MSTN-Trans. 0.54 0.15 5.07 4.80
MSTN-BiLSTM 4.39 0.81
Boubezoul [BOUBEZOUL2020105577] MSTN-Trans. 0.96 0.24 4.20 4.68
MSTN-BiLSTM 4.16 2.79
UCI-HAR [har240] MSTN-Trans. 29.98 3.48 45.30 4.36
MSTN-BiLSTM 7.14 2.03
PAMAP2 [pamap2_physical_activity_monitoring_231] MSTN-Trans. 1.58 0.20 0.46 3.21
MSTN-BiLSTM 0.76 0.086
ActBeCalf [DISSANAYAKE2025111462] MSTN-Trans. 0.89 0.14 8.53 2.20
MSTN-BiLSTM 4.14 1.76
MetroPT3 [metropt3] MSTN-Trans. 1.30 0.19 0.09 0.19
MSTN-BiLSTM 0.57 0.042
NASA [NasaTurboFan] MSTN-Trans. 1.31 0.45 4.52 2.31
MSTN-BiLSTM 4.25 1.48
2.7.2 Comparative Analysis and Edge-AI Deployability
Recent efficient architectures demonstrate varying complexity profiles: linear models such as DLinear [zeng2022transformerseffectivetimeseries] achieve $\mathcal{O}(CL)$ complexity, while spectral approaches like FEDformer [zhou2022fedformer] and TimesNet [wu2023timesnettemporal2dvariationmodeling] achieve $\mathcal{O}(CL\log L)$ complexity. Following the notation in SOFTS [han2024softsefficientmultivariatetime], their model achieves $\mathcal{O}(CL+CH)$ complexity via stochastic pooling, while iTransformer ($\mathcal{O}(C^{2}+CL+CH)$), PatchTST ($\mathcal{O}(CL^{2}+CH)$), and Standard Transformers ($\mathcal{O}(N\cdot CL^{2})$) incur heavier costs.
Against this landscape, the proposed MSTN-BiLSTM achieves strict linear temporal complexity $\mathcal{O}(CL+C^{2}+CK)$. Theoretically, this matches the optimal efficiency of SOFTS in the predominant real-world scenario where sequence length exceeds channel dimension ($L\gg C$). While SOFTS retains a theoretical advantage in high-dimensional, short-sequence settings ($C\gg L$) due to its purely linear channel dependence, such data structures are rare in long-term time series forecasting tasks. Meanwhile, MSTN-Transformer scaling as $\mathcal{O}(CL^{2}+C^{2}+CK)$ optimizes practical runtime by confining the quadratic term to a single encoder layer rather than a deep stack ($N$ layers) as seen in iTransformer or standard Transformer architectures. This strategic design allows MSTN to balance the high-capacity modeling of Transformers with the latency requirements of Edge-AI, delivering state-of-the-art performance with minimized computational overhead.
One thing to note is that despite the MSTN-BiLSTM’s better theoretical scaling ($\mathcal{O}(L)$) than the MSTN-Transformer’s $\mathcal{O}(L^{2})$, the latter runs significantly faster than the former (Table 5). This paradox is due to the fundamental conflict between the model’s structure and the way modern acceleration architectures process the data. The BiLSTM encoder is inherently recurrent, meaning that the calculation of the hidden state $h_{t}$ is strictly dependent on the previous state $h_{t-1}$. This sequential data dependency prevents the hardware’s parallel processing units from being fully utilized, forcing the computation into a slow, iterative loop (low hardware utilization). In contrast, the Transformer executes its $\mathcal{O}(L^{2})$ self-attention via highly optimized [vaswani2023attentionneed], complete matrix parallelization. The core operation, the $QK^{T}$ matrix multiplication, is performed simultaneously across thousands of processing elements. This high utilization of the hardware’s parallel capacity, confined to a single encoder layer before pooling, ensures the Transformer’s execution speed drastically outweighs the theoretical linearity of the BiLSTM’s sequential cost, resulting in lower practical latency across high-throughput platforms.
Standard Transformers maintain the full sequence length $L$ across a multi-layer stack of $N$ blocks for self-attention operations, resulting in cumulative attention cost proportional to $N\cdot\mathcal{O}(CL^{2})$.  In contrast, MSTN applies attention in its encoder, then immediately pools ($L\to 1$). Standard Transformers maintain $L$ across all layers, while MSTN eliminates $L$-dependence after encoding. Consequently, the remainder of the network operates on static feature vectors with $\mathcal{O}(1)$ complexity relative to $L$ for operations involving sequence length. By eliminating the recursive application of attention across deep layers, MSTN reduces the sequence-dependent FLOPs significantly. While MSTN-BiLSTM avoids quadratic attention costs entirely by employing BiLSTM encoding with $\mathcal{O}(CL)$ complexity. When combined with the remaining pipeline components (CNN, refinement, and projection), the total complexity is $\mathcal{O}(CL+C^{2}+CK)$, which scales linearly with sequence length $L$.
Following the baseline results and inference time metrics reported in SOFTS [han2024softsefficientmultivariatetime] on the Traffic data set (lookback window $L=96$, horizon $H=720$), we compare our MSTN results with their metrics reported in Table 5. MSTN-Transformer demonstrates that architectural efficiency can surpass asymptotic predictions. MSTN-Transformer achieves $69\times$ faster inference than PatchTST (0.72 ms vs. 50.00 ms). Furthermore, it outperforms models with better theoretical scaling, showing speedups of $34.7\times$ over SOFTS and $305\times$ over TimesNet. This practical gain, irrespective of the asymptotic class, is primarily delivered by the early temporal pooling of the MSTN architecture ($L\to 1$), a strategic design choice that dramatically reduces the overall FLOPs compared to deep, unpooled attention stacks. This performance represents a $22.4\times$ accuracy improvement and $34.7\times$ speedup over the SOFTS baseline, and a significant speed and accuracy advantage over all other models detailed in Table 5. This improved efficiency enables substantially reduced memory usage and inference latency, positioning MSTN as a practical solution for deployment on constrained hardware platforms.
As summarized in Table 5, MSTN demonstrates consistently efficient inference across six widely used long-term forecasting benchmarks. Despite the variation in dataset scale and temporal structure, MSTN maintains a compact model size in the range of 2.5–3.7 MB and sustains sub-millisecond latency across all horizons, including the longer 336 and 720-step settings. This stability in runtime behavior across heterogeneous datasets such as Traffic, ETTh2, ETTm2, Weather, Exchange, and ILI highlights the robustness of MSTN’s dual-path, multi-scale feature extraction. The uniformly low inference cost emphasizes its suitability for real-time forecasting scenarios and deployment in resource-constrained environments where computational consistency is critical.
@@ 2.7.2 Comparative Analysis and Edge-AI Deployability
3 Discussion
This work introduces MSTN, a DL framework founded on a hierarchical multi-scale modeling principle and defined by its strategic use of ETA. MSTN integrates core components: a multi-scale convolutional encoder that constructs a hierarchical feature pyramid; a sequential modeling component for long-term global dependency modeling, empirically validated with BiLSTM and Transformer variants; and a SGF mechanism augmented with SE and MHA to enable dynamic, context-aware feature integration. Existing methods often rely on fixed-scale structural priors (e.g., patch-based tokenization, fixed frequency transformations, or frozen backbone architectures). These limitations have historically forced a compromise between scale diversity and inference speed. In contrast, our approach efficiently integrates convolutional local pattern extraction with long-term global dependency modeling, leveraging ETA via an adaptive SGF mechanism. This enables our model to capture multi-scale temporal features while maintaining computational efficiency and simultaneously achieving SOTA performance.
@@ 3 Discussion
Although MSTN excels in standard forecasting regimes ($L\gg C$), we acknowledge that factorization-based models such as SOFTS retain a theoretical efficiency advantage in high-dimensional, short-sequence settings ($C\gg L$). However, since such datasets are rarely encountered in long-term time series forecasting tasks, this limitation has a negligible impact on practical deployability.
Future work will explore cross-variate attention mechanisms and large-scale pre-training to further enhance the capabilities of MSTN as a foundation model for time series understanding. Our work paves the way for efficient multi-scale temporal modeling under resource constraints, enabling real-time applications across diverse domains from healthcare to industrial monitoring.